OpenStudy (anonymous):

Using the compound interest formula A=P (1+ (r)/(n)^nt , find A when P=$1000, r=6%, t=5 years, and n=4

sam (.sam.):

A(5) = 1000(1 + 0.06/4)^(4*5) A(5) = 1000(1.015)^20 A(5) = 1000(1.346855007) A(5) = $1346.86